OpenStudy (anonymous):

if a function f is continuous for all real numbers and f(x) = (x^2-7x+12)/(x-4) when x doesnt equal 4, then f(4)=

OpenStudy (anonymous):

Factor \[f(x) = ((x -3) (x - 4))/ (x-4)\] f(4) = (4) - 3 f(4) = 1